# Words as Trigger Points in Social Media Discussions—A Large-Scale Case Study about UK Politics on Reddit

### v2

@@ Abstract.
Trigger points are a concept introduced by Mau et al. (2023) to study qualitative focus group interviews and understand polarisation in Germany. When people communicate, trigger points represent moments when individuals feel that their understanding of what is fair, normal, or appropriate in society is questioned. In the original studies, individuals show strong and negative emotional responses when certain triggering words or topics are mentioned. In this paper, we introduce the first systematic study of the large-scale effect of individual words as trigger points by analyzing a large amount of social media posts. We examine online deliberations on Reddit between 2020 and 2022 and collect >>>100 million posts from subreddits related to a set of words identified as trigger points in UK politics. We find that such trigger words affect user engagement and have noticeable consequences how online discussions occur. Trigger words cause animosity and incentivise hate speech, adversarial debates, and disagreements. Our work is the first to introduce trigger points to computational studies of online communication. The findings are relevant to researchers interested in affective computing, online deliberation, and those examining how citizens debate politics and society in light of affective polarisation.
Affective Computing, Polarization, Reddit, Trigger Words, User Engagement
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: THE WEB CONFERENCE 2025; 28 April–2 May, 2025; Sydney, Australiaccs: Information systems World Wide Webccs: Computing methodologies Natural language processing
1. Introduction
@@
Trigger points, introduced by Mau et al. (2023), are rooted in theories of affective political identity and relate to deeply lying beliefs about moral expectations and social dispositions. Examining trigger points in online discussions helps understand why and when social media users engage in disagreements or affective political deliberations. This opens the door to modelling social media user engagement more effectively and studying the conditions and causal mechanisms that lead to adverse reactions, hate speech, and abusive language in online debates.
@@
Figure 1. Consequences of trigger points. H1 refers to hypothesis 1, i.e., trigger words cause more engagement. H2 refers to hypothesis 2, i.e., trigger words cause more animosity.
@@
Trigger points are related to how individuals perceive themselves in society. People have subjective expectations about acceptable behaviour, such as explicit knowledge of law-like rules and, more importantly, an implicit understanding of appropriate actions. These deeply running beliefs of what is fair and unfair constitute a moral compass fundamental to navigating the social world. People feel triggered if these normative understandings are questioned. More formally, Mau et al. (2023, 246) define trigger points as “those moments that question what is acceptable against the backdrop of individuals’ understanding of the social contract. Trigger points jeopardise what individuals consider as the fabric of society and their role in it.” Trigger points can affect individuals through four mechanisms.
@@
  • Slippery slope. There could be a fear of society developing norms about appropriate behaviour that goes in subjectively undesired directions and the impression that the individual might not have the power to do something against it. “Opening the borders so that a country could be flooded with immigrants” is an example in this context.
@@
Going against deeply held beliefs about individuals and their societal role causes discomfort. People respond to such triggers in an affective behavioural mode and increase a debate’s intensity and emotionality.
@@
  • Hypothesis 1 more user engagement, which incites a higher number of messages or comments;
  • Hypothesis 2 higher levels of animosity, which incite polarisation, negativity, anger, and hate.
Our study focuses on how individual words can act as trigger points in online debates. We look into five words (Rwanda, Brexit, NHS, Vaccine, Feminism) we expect to work as trigger points in the context of UK politics. Collecting >>>100 million related Reddit posts, we use text analytics and NLP methods to understand whether these words lead to potentially negative, emotional, and harmful responses. The analysis builds upon a systematic comparison between a treatment and control group in three different scenarios:
  1. (1)
@@
    Semantics: words similar in meaning to the target words but without the triggering component.
  3. (3)
@@
    Time: time periods identified as relevant for each specific trigger word.
@@
Our results show that the trigger words cause higher engagement and more animosity, measured as an increase in a thread’s controversiality, negative sentiment, anger, and hate speech.
2. Related Work
@@
Trigger points (Mau et al., 2023) bear great potential to understand polarisation and division in online discussions. However, the existence or validity of the concept has not been tested on a large scale. Trigger points can be particularly relevant to work on social media. They help understand the causes of disagreements in online conversations (De Kock and Vlachos, 2021) and reasons for their escalation (De Kock and Vlachos, 2022; van der Meer et al., 2023). Further to triggering ad hominem attacks, escalation can lead to hate speech. There is, however, no unanimous agreement on how it differs from abusive and offensive language (Basile et al., 2019; Davidson et al., 2017; Talat and Hovy, 2016).
Despite the rising interest in hate speech and harmful language in social media platforms, including its implicit forms (Ousidhoum et al., 2019; Sap et al., 2019; ElSherief et al., 2021; Ocampo et al., 2023), and dog-whistles which are coded expressions that carry two meanings with a hateful or provocative coded one (Breitholtz and Cooper, 2021), we know little about the sources of these phenomena. Examining what can trigger harmful reactions may help us make more informed choices when collecting data, thus, leading to more robust detection models and mitigating the well-known problem of selection bias (Wiegand et al., 2019; Ousidhoum et al., 2020).
@@
Similarly, considering the broad array of harmful online behaviours, the wide range of elements they involve, and how they can be included in social media data (Olteanu et al., 2019), Blodgett et al. (2020) report a missing paradigm that helps assess the initial causes without mainly focusing on the performance of the detection model. Studying trigger points helps track the origins of bias and assess whether a social media post can lead to hateful comments (Dahiya et al., 2021).
Further, understanding how trigger words affect online communication also sheds light on how language affects polarisation on social media (Karjus and Cuskley, 2024). It lays the groundwork for understanding how polarised rhetoric may emerge and lead to intergroup conflict (Simchon et al., 2022).
3. Methodology
@@
We use the UK context as a large-scale case study to observe polarisation when individuals communicate on social media. Our goal is to report behavioural mechanisms similar to those observed by Mau et al. (2023) when conducting interviews to analyse the perceived polarisation of the German society. Based on a curated set of likely candidates, we test whether individual words can trigger the reactions outlined in the two hypotheses, i.e., (1) more engagement and (2) animosity.
3.1. Term selection
@@
We selected terms likely to provoke reactions from individuals in the United Kingdom. Specifically, we include National Health Service (NHS), Brexit, Rwanda, Vaccine, and Feminist/Feminism in our study. These trigger words were chosen because they are associated with high-profile public debates, policy changes, and social movements, which have been known to provoke significant public engagement and emotional responses. Table 1 indicates the respective mechanism through which the terms may trigger individuals.
@@
Table 1. Trigger words and their respective trigger mechanisms (described in the introduction).
@@
Rwanda plays a particular role in the UK’s immigration policy. [^1] https://www.gov.uk/government/publications/uk-rwanda-treaty-provision-of-an-asylum-partnership/uk-rwanda-treaty-provision-of-an-asylum-partnership-accessible Until late 2021, Rwanda was just the name of any other African country. However, in early 2022 the term’s meaning became emblematic of how the British government suggested handling asylum seekers, and the home secretary struck a deal with Rwanda: the Migration and Economic Development Partnership. Instead of processing asylum cases directly in the UK, migrants would be flown to Kigali, and their claims were to be assessed there, in turn for U\$300m, (2.3% of Rwandan GDP) and expenses for every migrant sent. With this, Rwanda became a trigger word for discussions on immigration policy in the UK from 2022 onwards, particularly related to “unfair favours towards migrants and unregulated immigration” (The Economist, 2023).
@@
The UK’s departure from the European Union has been controversial since its inception. Even eight years post-exit, debates persist regarding the soundness of the decision. The division among the population was already evident during the tightly contested 2016 referendum, which revealed distinct trends among various demographic groups based on location, age, and income (Arnorsson and Zoega, 2018). Since the referendum, further divisions have emerged concerning the future relationship between the UK and EU (Vasilopoulou and Talving, 2019). Additionally, misinformation and disinformation have intensified these debates (Höller, 2021; Bruno et al., 2022). Here, considerations about treating EU migrants differently than the national population (inequality), norm transgressions of “criminal” migrants, particularly from the East of Europe, or opening the gate to “unlimited” immigration make Brexit a potential trigger word for analysis.
@@
Due to financial austerity measures, the UK government was accused of reducing its budget and neglecting the health care system (The Guardian, 2022; The Independent, 2024). As a result, the NHS has become a hot topic of discussion among the UK population. Public opinion is divided, with some supporting that it receives adequate funding, while others believe it wastes resources or that its services are in decline (Gershlick et al., 2015). This trigger word is related to going against considerations about unequal treatment (i.e., inequality as shown in Table 1).
@@
Public vaccination programs, particularly those against COVID-19, have caused significant debate in recent years. Even after their widespread roll-out, discussions over their effectiveness and safety are ongoing. Initial reluctance was observed across various demographic segments, influenced by factors such as education level, political affiliation, and socioeconomic status (Troiano and Nardi, 2021). Post-roll-out, debates have shifted to focus on issues such as the fairness of vaccine distribution and unreasonable demands for vaccination (Nichol and Mermin-Bunnell, 2021). Moreover, the proliferation of misinformation and conspiracy theories has further fuelled public division (Hayawi et al., 2022; Garett and Young, 2021), particularly in social media, making Vaccine an ideal trigger word for our study.
Threads Comments Year
Rwanda 0.05 1.87 2020, 2022
Brexit 0.37 11.57 2020 - 2022
NHS 0.34 6.20 2020 - 2022
Vaccine 1.42 28.11 2019, 2022
Feminism 0.94 17.79 2020 - 2022
EU 0.2 12.4 2020 - 2022
Red Cross 0.02 0.2 2020 - 2022
SubSaharan 0.08 1.5 2020, 2022
Surgery 0.5 3.9 2019, 2022
Woman 2.1 19.3 2020 - 2022
Table 2. Numbers, in millions, of threads, comments, and time period for each trigger word.
@@
Misogyny can be observed on social media, where anonymity often allows users to express offensive and hateful sentiments (Barker and Jurasz, 2019). Previous work has shown that women politicians and journalists tend to be harassed online every few seconds (Posetti et al., 2022).  [^2] https://www.amnesty.org.uk/press-releases/women- abused-twitter-every-30-seconds-new-study Some social media influencers have built a large audience by sharing and promoting harmful content against women. This has led to the formation of online echo chambers (Farrell et al., 2019) where such views are amplified and propagated. Our research includes Feminism as a trigger word based on mechanisms related to norm transgressions, fears of a slippery slope and unreasonable behavioural demands.
@@
Table 3. Subreddits selected for each trigger word.
@@
Table 4. Percentage of comments per thread before and after each trigger appearance (Hypothesis 1). Difference-in-Difference (DiD) and 95% confidence interval (CI) included for the Control, Space and Semantics, comparisons. Bold scores indicate statistically significant results. Detailed distribution of comments can be seen in Appendix B, see Figure 7.
3.2. Data collection
Our study uses data from Reddit comments in English, collected via the Pushshift API (Baumgartner et al., 2020). We searched for exact matches of specific keywords in comments from 2019 to 2022, and then collected the entire discussion thread for each extracted comment [^3] When extracting comments for the term woman, we capped the maximum number of comments at 20 million due to the large volume.. We excluded posts in languages other than English, as identified with fastText-based language identifier (Bojanowski et al., 2017). Posts marked as “[deleted]” by the API were also removed. Following this, only threads that contained posts before and after the appearance of the target keywords were retained, as it is crucial to analyse the potential differences before and after the appearance of a given trigger point.
As a final filtering step we limited our analysis to posts occurring 30 minutes before and 30 minutes after the first mention of the triggering word. Previous research has shown that Reddit posts typically receive the majority of interactions within a short window after the post is created (Thukral et al., 2018). Moreover, early responses are often important in shaping the direction of the conversation (Singer et al., 2016; Horne et al., 2017). Considering the fast-paced nature of Reddit, the 30 minutes window filters out unrelated responses and ensures the focus of our analysis remains on the most relevant data. Table 2 shows the total number of comments, threads, and comments collected for each trigger word and the period they were gathered.
3.3. Research design
We identify the causal effect of trigger points with the difference-in-difference estimator (DiD) which systematically compares a treatment with a control group (Angrist and Pischke, 2009; Card and Krueger, 1994). It calculates the causal effect of an intervention as the difference before and after the intervention in the treatment group minus the difference before and after the intervention in the control group. In our context, we expect users in the treatment group to be highly likely to be triggered by the respective terms and those in the control group to be much less triggered by the respective words. There should be a more substantive behavioural change for the treatment group than for the control group.
3.3.1. Control groups

### v3

# Words as Trigger Points in Social Media Discussions: A Large-Scale Case Study about UK Politics on Reddit

@@ Abstract
Political debates on social media sometimes flare up. From that moment on, users engage much more with one another; their communication is also more emotional and polarised. While it has been difficult to grasp such moments with computational methods, we suggest that trigger points are a useful concept to understand and ultimately model such behaviour. Established in qualitative focus group interviews to understand political polarisation (Mau, Lux, and Westheuser 2023), trigger points represent moments when individuals feel that their understanding of what is fair, normal, or appropriate in society is questioned. In the original studies, individuals show strong and negative emotional responses when certain triggering words or topics are mentioned. Our paper finds that these trigger points also exist in online debates. We examine online deliberations on Reddit between 2020 and 2022 and collect $>$100 million comments from subreddits related to a set of words identified as trigger points in UK politics. Analysing the comments, we find that trigger words increase user engagement and animosity, i.e., more negativity, hate speech, and controversial comments. Introducing trigger points to computational studies of online communication, our findings are relevant to researchers interested in affective computing, online deliberation, and how citizens debate politics and society in light of affective polarisation.
1 Introduction
Online debates may suddenly escalate, and users intensely engage in emotional and confrontational deliberations. While these turning points are difficult to model and predict, they are essential moments in online communication. They mark the moment when a discussion is likely to get off the rails, and social media users stop debating with one another and instead debate against one another. Our study introduces trigger points as a useful concept to understand why and when social media users engage in such behaviour.
@@ 1 Introduction
Trigger points, introduced by Mau, Lux, and Westheuser (2023), are rooted in theories of affective political identity and relate to deeply lying beliefs about moral expectations and social dispositions. They open the door to modelling social media user engagement more effectively and studying the conditions and causal mechanisms that lead to adverse reactions, hate speech, and abusive language in online debates.
@@
Figure 1: Consequences of trigger points. H1 refers to hypothesis 1, i.e., trigger words cause more engagement. H2 refers to hypothesis 2, i.e., trigger words cause more animosity.
@@
Trigger points are related to how individuals perceive themselves in society. People have subjective expectations about acceptable behaviour, such as explicit knowledge of law-like rules and, more importantly, an implicit understanding of appropriate actions. These deeply rooted beliefs of what is fair and unfair constitute a moral compass fundamental to navigating the social world. People feel triggered if these normative understandings are questioned. More formally, Mau, Lux, and Westheuser (2023, 246) define trigger points as “those moments that question what is acceptable against the backdrop of individuals’ understanding of the social contract. Trigger points jeopardise what individuals consider the fabric of society and their own role in it.” The authors suggest that trigger points can affect individuals through four mechanisms.
@@
  • Slippery slope. There could be a fear of society developing norms about appropriate behaviour that goes in subjectively undesired directions and the impression that the individual might not have the power to do something against it. Claims such as “Opening the borders so that a country could be flooded with immigrants” is an example.
@@
Going against deeply held beliefs about individuals and their societal role causes discomfort. People respond to such triggers in an affective behavioural mode, increasing a debate’s intensity and emotionality.
@@
  • Hypothesis 1: more user engagement, which leads to more messages or comments.
  • Hypothesis 2: higher levels of animosity, which causes polarisation, negativity, anger, and hate.
Our study seeks to establish the concept of trigger points in online communication. The research design and case selection closely follow this logic and examine whether and how individual words can act as trigger points in online debates. To this end, we conduct a large-scale case study and select five words that are ex-ante likely to work as trigger points in the context of UK politics. Specifically, we look into the words Rwanda, Brexit, NHS, Vaccine, Feminism. Collecting $>$100 million related Reddit comments, we use text analytics and NLP methods to understand whether these words lead to potentially negative, emotional, and harmful responses. The analysis builds upon a systematic comparison between a treatment and control group in three different scenarios:
  1. 1.
@@
    Contextual: words that appear in a similar context to the target words but without the triggering component.
  3. 3.
@@
    Time: periods identified as relevant for each specific trigger word.
@@
We find that the trigger words cause higher engagement and more animosity, measured as an increase in a thread’s controversiality, negative sentiment, anger, and hate speech.
2 Related Work
Existing work studies how particular political words or phrases lead to more engagement (Brady et al. 2017; Hameleers et al. 2021) and posts with more emotion and animosity (Friedman et al. 2004; Martin et al. 2021; Schroth, Bain-Chekal, and Caldwell 2005; Rho and Mazmanian 2020) in various online settings. However, they lack a joint conceptual and theoretical framework—something we aim to provide.
@@ 2 Related Work
Trigger points are fundamental to understanding polarisation and division in online discussions. Given that they are at the root of different conflictive behaviours later on, they help understand the causes of disagreements in online conversations (De Kock and Vlachos 2021) and their escalation (De Kock and Vlachos 2022; van der Meer et al. 2023) triggering explicit and implicit ad hominem attacks, hate speech, abusive or offensive language on social media (Talat and Hovy 2016; Davidson et al. 2017; Basile et al. 2019; Ousidhoum et al. 2019; Sap et al. 2019; ElSherief et al. 2021; Ocampo et al. 2023).
@@
Similarly, considering the broad array of harmful online behaviours, the wide range of elements they involve, and how they can be included in social media data (Olteanu et al. 2019), Blodgett et al. (2020) report a missing paradigm that helps assess the initial causes without mainly focusing on the performance of the detection model. Studying trigger points helps track the origins of bias, assess whether a social media post can lead to toxic comments (Dahiya et al. 2021), and how language affects polarisation on social media (Karjus and Cuskley 2024). It lays the groundwork for understanding how polarised rhetoric may emerge and lead to intergroup conflict (Simchon, Brady, and Van Bavel 2022).
Related to trigger points are dog-whistles—coded expressions that carry two meanings: a seemingly harmless one, and another one with a hidden hateful or provocative connotation that only insiders can decipher (Breitholtz and Cooper 2021). In social sciences, the concept of dog whistle has its roots in the study of racial politics in the United States, leading politicians to develop more subtle ways of appealing to racial prejudices (López 2013) and has evolved to encompass a broader range of phenomena, ranging from perpetuated systemic racism (Filimon and and 2024), to more broadly communication strategies of political ambiguity in identity politics(Tolvanen 2024), such as religion (Perry 2023) or immigration (Filimon 2016). Trigger points are similar in their equivocality and potential to evoke emotional responses. Conversely, using a dog whistle is conscious, but there is typically no intent when relying on trigger points. Moreover, while dog whistles are rhetorical devices to hide the real intention from some and create a rally-around-the-flag effect, trigger points are not hidden communication and incite a strong sense of otherness towards those covered by the trigger.
Research on the automation of identifying affective triggers in online discussions remains limited. For instance, Almerekhi et al. (2022) developed a framework to identify triggers for toxic behaviour within Reddit discussions. Similar work typically uses a restricted set of lexical features, primarily sentiment analysis, and often examines the level of engagement some topics may generate (Aldous, An, and Jansen 2023). Others, such as Beel et al. (2022), explicitly investigate highly controversial topics like abortion, climate change, and gun control, attempting to predict contentiousness within discussions, often heavily relying on Reddit comment upvote ratios as a proxy for contentiousness. Finally, some studies focus on online conversational tensions by predicting conversational derailment, considering the overall progression of a conversation rather than specific terms as causes (Zhang et al. 2018; Hessel and Lee 2019).
3 Methodology
@@
To empirically demonstrate that individual words can trigger online debates, we look into most likely cases in the context of UK politics. Based on a curated set of likely candidate words for the political context in the UK (see Section 3.1), we test whether individual words can trigger the reactions outlined in the two hypotheses, i.e., (1) more engagement and (2) animosity. Our goal is to report behavioural mechanisms similar to those observed by Mau, Lux, and Westheuser (2023) when conducting interviews to analyse the perceived polarisation. If we could not find these effects in these intuitive cases, we would falsify the theory for online contexts.
3.1 Term selection
@@
We selected National Health Service (NHS), Brexit, Rwanda, Vaccine, and Feminist/Feminism. These terms are associated with high-profile public debates, policy changes, and social movements in the UK. Hence, they typically provoke significant public engagement and emotional responses. We expect them to cause the intended reactions in British social media users. Table 1 indicates mechanisms through which the terms may trigger individuals.
@@
Table 1: Trigger words and their respective trigger mechanisms (described in the introduction).
@@
Until late 2021, Rwanda was just the name of any other African country. However, in early 2022, the term’s meaning became emblematic of how the British government suggested handling asylum seekers, and the home secretary struck a deal with Rwanda: the Migration and Economic Development Partnership. [^1] https://www.gov.uk/government/publications/uk-rwanda-treaty-provision-of-an-asylum-partnership/uk-rwanda-treaty-provision-of-an-asylum-partnership-accessible With this agreement, Rwanda became a trigger word for discussions on immigration policy in the UK from 2022 onwards, particularly related to “unfair favours towards migrants and unregulated immigration” (The Economist 2023).
@@
The UK’s departure from the European Union has been controversial since its inception. The division among the population was already evident during the tightly contested 2016 referendum(Arnorsson and Zoega 2018). Additionally, misinformation and disinformation have intensified these debates (Höller 2021; Bruno, Lambiotte, and Saracco 2022). Here, considerations about treating EU migrants differently than the national population (inequality), norm transgressions of “criminal” migrants, particularly from Eastern Europe, or opening the gate to “unlimited” immigration make Brexit a potential trigger word for analysis.
@@
Due to financial austerity measures, the UK government was accused of reducing its budget and neglecting the healthcare system (The Guardian 2022; The Independent 2024). Public opinion is divided, with some supporting that the NHS receives adequate funding, while others believe it wastes resources or that its services are in decline (Gershlick, Charlesworth, and Taylor 2015). This trigger word is related to going against considerations about unequal treatment (i.e., inequality as shown in Table 1).
@@
Public vaccination programs, particularly those against COVID-19, have caused significant debate in recent years. Post-roll-out, debates have shifted to focus on issues such as the fairness of vaccine distribution and unreasonable demands for vaccination (Nichol and Mermin-Bunnell 2021). Moreover, the proliferation of misinformation and conspiracy theories has further fuelled public division (Hayawi et al. 2022; Garett and Young 2021), particularly in social media, making Vaccine an ideal trigger word for our study.
@@
Misogyny can be observed on social media, where anonymity often allows users to express offensive and hateful sentiments (Barker and Jurasz 2019). Some social media influencers have built a large audience by sharing and promoting harmful content against women. This has led to the formation of online echo chambers (Farrell et al. 2019) where such views are amplified and propagated. Our research includes Feminism as a trigger word based on mechanisms related to norm transgressions, fears of a slippery slope, and unreasonable behavioural demands.
@@
Table 2: Subreddits selected for each trigger word.
@@
Table 3: Percentage of comments per thread before (Be) and after (Af) each trigger appearance (Hypothesis 1). Difference-in-Difference (DiD) and 95% confidence interval (CI) included for the Control, Space and Contextual, comparisons. Bold scores indicate statistically significant results. Detailed distribution of comments can be seen in Appendix B, see Figure 8.
3.2 Data collection
Our study uses data from Reddit comments in English, collected via the Pushshift API (Baumgartner et al. 2020). We searched for exact matches of specific keywords in comments from 2019 to 2022 and then collected the entire discussion thread for each extracted comment. [^2] When extracting comments for the term woman, we capped the maximum number of comments at 20 million due to the large volume. We excluded comments in languages other than English, identified as such using a fastText-based language identifier (Bojanowski et al. 2017). Comments marked as “[deleted]” by the API were also removed. Following this, only threads that contained comments before and after the appearance of the target keywords were retained, as it is crucial to analyse the potential differences before and after the appearance of a given trigger point.
As a final filtering step, we limited our analysis to comments 30 minutes before and 30 minutes after the first mention of the triggering word. Previous research has shown that Reddit comments typically receive the majority of interactions within a short window after the comment is created (Thukral et al. 2018). Moreover, early responses are often important in shaping the direction of the conversation (Singer et al. 2016; Horne, Adali, and Sikdar 2017). Considering the fast-paced nature of Reddit, the 30-minute window filters out unrelated responses and ensures the focus of our analysis remains on the most relevant data [^3] Table 7 in Appendix B shows the total number of comments, threads, and comments collected for each trigger word and the period they were gathered..
3.3 Research design
We identify the causal effect of trigger points with the difference-in-difference estimator (DiD), which systematically compares a treatment with a control group (Angrist and Pischke 2009; Card and Krueger 1994). It calculates the causal effect of an intervention as the difference before and after the intervention in the treatment group minus the difference before and after the intervention in the control group. In our context, we expect users in the treatment group to be highly likely to be triggered by the respective terms and those in the control group to be much less triggered by the respective words. There should be a more substantive behavioural change for the treatment group than for the control group.
Control groups